# Learning-to-Defer with Expert-Conditioned Advice

### 15 Mar 2026 v1

@@ Abstract
Learning-to-Defer routes each input to the expert that minimizes expected cost, but it assumes that the information available to every expert is fixed at decision time. Many modern systems violate this assumption: after selecting an expert, one may also choose what additional information that expert should receive, such as retrieved documents, tool outputs, or escalation context. We study this problem and call it Learning-to-Defer with advice. We show that a broad family of natural separated surrogates, which learn routing and advice with distinct heads, are inconsistent even in the smallest non-trivial setting. We then introduce an augmented surrogate that operates on the composite expert–advice action space and prove an $\mathcal{H}$-consistency guarantee together with an excess-risk transfer bound, yielding recovery of the Bayes-optimal policy in the limit. Experiments on tabular, LLMs, and multi-modal tasks show that the resulting method improves over standard Learning-to-Defer while adapting its advice-acquisition behavior to the cost regime.
@@ 1 Introduction
Modern decision systems rarely act on a fixed snapshot of information. A retrieval-augmented model may be chosen to generate an answer and then be provided with specific retrieved documents (Guu et al., 2020; Lewis et al., 2020). A tool-augmented agent may be assigned a task and then granted access to a calculator, a search engine, or a code interpreter (Schick et al., 2023). In each of these settings, two decisions are made in sequence: who should handle the instance, and what additional information, if any, should be acquired to help them. The value of that additional information is not known when the decision is made; it is revealed only after the commitment. What, then, should the learner optimize: the sequential control flow, or the final executed decision?
Learning-to-Defer provides a principled framework for the first of these two decisions. Given an input $x$, the system routes each instance to the expert best suited to handle it, balancing predictive performance against expert-specific costs (Madras et al., 2018; Mozannar and Sontag, 2020; Mao et al., 2023a; Montreuil et al., 2026a). This framework has produced strong theoretical foundations and practical successes. Yet it operates under a fundamental assumption: the information available to every expert is fixed at decision time. The system chooses who acts, but never what they get to see. The entire question of whether and what to acquire is outside the model. This misses an important class of systems in which the value of extra information depends on the expert who will consume it, and can even change which expert is optimal. Recent work has begun to study the binary ask/not-ask case of querying human feedback for a single enriched predictor (Pugnana et al., 2025); our focus here is the more general routed setting in which the value of advice depends on the expert that will use it. In this richer setting, the central statistical question is Bayes consistency: can one learn a policy that recovers the Bayes-optimal executed expert-advice decision, or equivalently the best expert under its best advice choice?
This coupling makes the problem qualitatively different from standard learning to defer. A natural response is to keep the sequential structure and learn routing and advice with separate heads. But does such a decomposition preserve the Bayes rule of the underlying protocol? We show that, in general, it does not. Even in the simplest setting — two experts and a single binary advice — the separated surrogate is inconsistent for the true loss. The issue is not optimization, but statistical alignment: the protocol executes an expert–advice pair, while the separated surrogate compares the wrong objects.
We call this framework Learning-to-Defer with advice. It generalizes classical Learning-to-Defer by introducing expert-conditioned advice acquisition as a first-class component of the deferral decision. Can one still obtain a learning problem that is both statistically correct and simple to optimize, despite the negative result above? We show that the answer is yes. We introduce a surrogate with an explicit Bayes-consistency guarantee: its minimization is provably aligned with the Bayes-optimal executed expert-advice decision of the original protocol, that is, with the expert whose best advice yields the smallest conditional expected cost.
@@
(i) we introduce Learning-to-Defer with advice, a formulation in which the system jointly selects an expert and an expert-conditional advice action, and we derive the corresponding Bayes-optimal policy, which compares experts at their best-advised cost (Section 4, especially Section 4.3 and Lemma 3); (ii) we prove that a broad family of natural separated router-query surrogates can be Fisher inconsistent for the true protocol loss, even in the simplest binary-advice setting (Section 4.4 and Theorem 6); (iii) we introduce an augmented surrogate and prove an $\mathcal{H}$-consistency guarantee together with an excess-risk transfer bound, thereby establishing Bayes consistency and showing that the learned policy recovers the Bayes-optimal expert-advice decision in the limit, equivalently the expert whose best advice yields the smallest conditional expected cost (Section 4.5, Lemma 7, Theorem 8, and Corollary 9); and (iv) we validate the framework on tabular, language, multi-modal, and synthetic experiments, where it improves over standard Learning-to-Defer, adapts its advice-acquisition rate to the cost regime, and empirically exhibits the failure mode predicted for separated surrogates (Section 5 and Appendix C.2).
@@ 3 Preliminaries
We consider a fixed collection of $J$ experts, indexed by $J]\coloneqq\{1,\dots,J\}$. Each expert $j\in[J]$ is associated with a measurable cost function $c_{j}:\mathcal{X}\times\mathcal{Y}\to[0,C]$, uniformly bounded by some $C<\infty$, that quantifies the expense of routing $(x,y)$ to expert $j$. This formulation is agnostic to the nature of the expert: an expert may be a predictive model, a language model system, a human annotator, or any other decision mechanism that induces measurable costs. We model expert prediction as $e_{j}:\mathcal{X}\to\mathcal{Y}$ and decompose the cost as $c_{j}(x,y)\;\coloneqq\;\psi\bigl(e_{j}(x),\,y\bigr)+\beta_{j}$, where $\psi:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a measurable task-specific loss (e.g., the 011 loss for classification) and $\beta_{j}\geq 0$ is an expert consultation cost (Madras et al., 2018; Mozannar and Sontag, 2020; Verma et al., 2022). We denote the tuple of deterministic expert predictors, when this special case is used, by $\mathbf{e}\coloneqq(e_{1},\dots,e_{J})$.
Given a collection of fixed, pre-trained experts, Learning-to-Defer seeks a routing policy that assigns each instance to the most cost-effective expert (Madras et al., 2018; Mozannar and Sontag, 2020; Narasimhan et al., 2022; Mao et al., 2023a, 2024c; Montreuil et al., 2025b, 2026a). Let $\mathcal{H}_{r}\subseteq\{s_{r}:\mathcal{X}\times[J]\to\mathbb{R}\}$ denote a hypothesis class of measurable scoring functions for routing. Each $s_{r}\in\mathcal{H}_{r}$ induces a measurable router $r:\mathcal{X}\to[J]$ via $r(x)=\operatorname*{arg\,max}_{j\in[J]}s_{r}(x,j)$; we write $\mathbf{s}_{r}(x)=(s_{r}(x,j))_{j\in[J]}$ for the full score vector. The deferral loss incurred by routing a realization $(x,y)$ through $r$ is
@@
The indicator functions in $\ell_{\mathrm{def}}$ make it discontinuous and non-differentiable, precluding direct gradient-based optimization (Zhang, 2002; Steinwart, 2007; Mozannar and Sontag, 2020). Standard practice replaces it with a convex surrogate $\Phi_{\mathrm{def}}$ that operates on the scoring function $s_{r}$ and is amenable to gradient-based methods. The surrogate deferral risk is $\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})=\mathbb{E}_{(X,Y)\sim\mathcal{D}}[\Phi_{\mathrm{def}}(s_{r};\,X,Y,\mathbf{e})]$, with best-in-class value $\mathcal{E}_{\Phi_{\mathrm{def}}}^{\ast}(\mathcal{H}_{r})\coloneqq\inf_{s_{r}\in\mathcal{H}_{r}}\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})$. To obtain population-level excess-risk control, Awasthi et al. (2022) introduced $\mathcal{H}$-consistency bounds. Applied to $\mathcal{H}_{r}$, the bound takes the following form.
@@
When $\mathcal{H}_{r}$ is rich enough for these gaps to vanish, the bound recovers Bayes consistency (Steinwart, 2007; Awasthi et al., 2022). In the standard L2D setting, such bounds have been established for the comp-sum family of cross-entropy surrogates (Mao et al., 2023b, 2024a), guaranteeing that surrogate minimization over a rich enough class recovers the Bayes-optimal router. In the next section, we extend the routing problem to include advice acquisition and examine which surrogates preserve consistency in this richer setting.
@@ 4.2 The Deferral-Advice Loss
Formally, a deferral-advice policy is a pair of measurable maps $r:\mathcal{X}\to[J]$ and $q:\mathcal{X}\times[J]\to[K]_{0}$. Let $\mathcal{H}_{q}\subseteq\{s_{q}:\mathcal{X}\times[J]\times[K]_{0}\to\mathbb{R}\}$ be a hypothesis class of measurable query scoring functions; each $s_{q}\in\mathcal{H}_{q}$ induces a query function via $q(x,j)=\operatorname*{arg\,max}_{k\in[K]_{0}}s_{q}(x,j,k)$. For a fixed input $x$, let $\mathbf{S}_{q}(x)\in\mathbb{R}^{J\times(K+1)}$ denote the query-score matrix with entries $s_{q}(x,j,k)$. Equivalently, taking the argmax along each row yields one candidate advice index for each expert. Upon observing $x$, the system routes to expert $j=r(x)$ and selects advice index $k=q(x,j)$; the masked advice $\widetilde{A}^{(k)}$ is then revealed to expert $j$. The query function $q(x,\cdot)$ is defined for every expert, but only the executed value $q(x,r(x))$ enters the realized loss: the remaining values $q(x,j^{\prime})$ for $j^{\prime}\neq r(x)$ are counterfactual and invisible to the protocol. Appendix Algorithm 2 records this inference rule explicitly.
@@
The double indicator selects exactly one expert–advice pair per instance. The loss has the same structure as the standard deferral loss (1) but ranges over $J(K+1)$ actions rather than $J$, and it depends on the policy only through the executed pair (r(x),q(x,r(x)))(r(x),\,q(x,r(x))). When only the null advice action is available, the query function is trivial, $c_{j,0}(x,a,y)=c_{j}(x,y)$, and the loss reduces to $\ell_{\mathrm{def}}$ of (1); thus the standard deferral setting is a special case (Mozannar and Sontag, 2020; Mao et al., 2023a; Montreuil et al., 2026a).
@@ 4.3 Deferral with Advice Dominates Standard Deferral
The proof is given in Appendix B.1. The loss depends on the policy only through the executed pair (r(x),q(x,r(x)))(r(x),\,q(x,r(x))), so the population risk can be minimized pointwise in $x$ over the finite cost table. The result reveals a fundamental coupling: the router does not compare experts on their unaided costs $\mathbb{E}[c_{j}(X,Y)\mid X=x]$ as in the standard setting, but on their optimally-advised ones. An expert that would never be selected under standard deferral may become optimal once paired with the right advice. The two decisions—who acts and what they see—cannot be separated without loss.
@@
For every expert $j\in[J]$ and $\mathbb{P}_{X}$-a.e. $x$,
Consequently, if $\mathcal{E}^{\star}_{\ell_{\mathrm{def}}}$ denotes the Bayes risk of the standard deferral setting of Section 3, then
@@
Equality holds if and only if, for $\mathbb{P}_{X}$-a.e. $x$,
$\min_{j\in[J]}\min_{k\in[K]_{0}}\mathbb{E}\bigl[c_{j,k}(X,A,Y)\mid X=x\bigr]=\min_{j\in[J]}\mathbb{E}\bigl[c_{j,0}(X,A,Y)\mid X=x\bigr].$
@@ 4.4 The Separated Surrogate Fails
Because the protocol is sequential, the natural first attempt is to learn it sequentially as well: one router score, one query head per expert, and smooth logistic replacements for the hard indicators (Bartlett et al., 2006). But does mirroring the control flow suffice for consistency? It does not. We show that a broad family of separated surrogates already fails in the smallest non-trivial case: two experts $(J=2)$ and one binary advice $(K=1)$.
@@
has Bayes-optimal expert $j^{\star}=1$, while the unique minimizer of the pointwise surrogate (10) decodes to expert $2$. In particular, the entire family (10) is not Fisher consistent for the true protocol loss.
@@ 4.5 The Augmented Surrogate
This reformulation does not alter the underlying decision problem. A sequential policy $(r,q)$ induces the composite action $\pi(x)=(r(x),\,q(x,r(x)))$, and every composite action $(j,k)$ can in turn be implemented by a sequential policy with the same realized loss. The composite view simply makes explicit what the loss has depended on all along: the executed expert–advice pair. Proposition 16, proved in Appendix B.5, shows that the sequential and composite formulations therefore induce the same population risks and share the same Bayes-optimal decisions as those identified in Lemma 3.
@@
The composite view aligns the learning problem with the Bayes comparison itself. One still faces the non-differentiability of the true loss from Definition 2, so the remaining task is to construct a surrogate on the augmented action space. The key step, proved in Appendix B.6, is a decomposition of the realized loss into an action-independent offset and a weighted sum of composite mismatch indicators $\mathbf{1}\{\pi(x)\neq(j,k)\}$. Since the offset does not affect the pointwise comparison across actions, surrogate design can focus entirely on the mismatch term. The resulting problem is a weighted multiclass prediction task over $\Pi$.
@@
We now establish the central theoretical guarantee of the paper: the augmented surrogate does not merely avoid inconsistency, it admits a quantitative excess-risk transfer bound for the entire comp-sum family and hence for every $\tau\geq 0$.
@@
Corollary 9 (Asymptotic Bayes-risk consistency).
@@
Corollary 9 is the statistical payoff: whenever the hypothesis class is rich enough and the surrogate is well-optimized, the induced policy converges to the Bayes-optimal deferral-advice risk (Steinwart, 2007). When $\mathcal{H}_{\pi}$ contains all measurable functions, both minimizability gaps vanish by definition (Zhang and Agarwal, 2020; Bartlett et al., 2006; Long and Servedio, 2013), and the bound reduces to classical Bayes-consistency. The separated surrogate offers no such guarantee—its minimizers can converge to a provably suboptimal policy regardless of model capacity.

### 19 Mar 2026 v2

@@ Abstract
Learning-to-Defer routes each input to the expert that minimizes expected cost, but it assumes that the information available to every expert is fixed at decision time. Many modern systems violate this assumption: after selecting an expert, one may also choose what additional information that expert should receive, such as retrieved documents, tool outputs, or escalation context. We study this problem and call it Learning-to-Defer with advice. We show that a broad family of natural separated surrogates, which learn routing and advice with distinct heads, is inconsistent even in the smallest non-trivial setting. We then introduce an augmented surrogate that operates on the composite expert–advice action space and prove an $\mathcal{H}$-consistency guarantee together with an excess-risk transfer bound, yielding recovery of the Bayes-optimal policy in the limit. Experiments on tabular, language, and multi-modal tasks show that the resulting method improves over standard Learning-to-Defer while adapting its advice-acquisition behavior to the cost regime; a synthetic benchmark confirms the failure mode predicted for separated surrogates.
@@ 1 Introduction
Modern decision systems face two intertwined choices: who should handle a given instance, and what information that expert should receive. A retrieval-augmented model may answer only after receiving retrieved documents (Guu et al., 2020; Lewis et al., 2020). A tool-augmented agent may be granted access to a calculator, a search engine, or a code interpreter (Schick et al., 2023). These examples concern the second choice; standard Learning-to-Defer (Madras et al., 2018; Mozannar and Sontag, 2020; Mao et al., 2023a; Montreuil et al., 2026a), by contrast, addresses the first. Our setting combines both: the learner selects an expert and, conditional on that choice, decides what advice to acquire — before learning whether that advice will help. What, then, should the learner optimize: the sequential control flow, or the final executed decision?
Learning-to-Defer (L2D) treats the information available to every expert as fixed at decision time. The entire question of whether and what to acquire is outside the model. This misses an important class of systems in which the value of extra information depends on the expert who will consume it, and can even change which expert is optimal. While recent work studies the binary ask/not-ask case for a single enriched predictor (Pugnana et al., 2025), our focus is the more general routed setting in which advice is expert-conditioned and may come from multiple sources. We call this setting Learning-to-Defer with advice: the system jointly selects an expert and decides what information that expert should receive. The central statistical question is then Bayes consistency (Zhang, 2002; Steinwart, 2007; Bartlett et al., 2006): can one learn a policy that recovers the Bayes-optimal expert-advice decision, that is, the expert whose best advice yields the smallest conditional expected cost?
Because routing and advice are coupled — the best advice depends on the expert, and the best expert depends on the advice — one might keep the sequential structure and learn them with separate heads. We show that this natural decomposition is, in general, inconsistent: even with two experts and binary advice, the separated surrogate fails to recover the Bayes rule. The key insight is that the protocol evaluates expert–advice pairs, so the surrogate must operate on the same composite action space. Building on this observation, we introduce an augmented surrogate over the joint expert–advice space and prove an $\mathcal{H}$-consistency guarantee: its minimization is provably aligned with the Bayes-optimal expert–advice decision.
@@
  • We introduce Learning-to-Defer with advice, a formulation in which the system jointly selects an expert and an expert-conditional advice action, and derive the Bayes-optimal policy, which compares experts at their best-advised cost (Section 4.3, Lemma 3).
  • We prove that a broad family of natural separated router-query surrogates is not Bayes consistent, even in the simplest binary-advice setting (Section 4.4, Theorem 6).
  • We introduce an augmented surrogate and prove an $\mathcal{H}$-consistency guarantee together with an excess-risk transfer bound, establishing that surrogate minimization recovers the Bayes-optimal expert-advice decision in the limit (Section 4.5, Theorem 8, Corollary 9).
  • We validate the framework on tabular, language, multi-modal, and synthetic experiments, where it improves over standard Learning-to-Defer, adapts its advice-acquisition rate to the cost regime, and empirically exhibits the failure mode predicted for separated surrogates (Section 5, Appendix C.2).
@@ 3 Preliminaries
We consider a fixed collection of $J$ experts, indexed by $J]\coloneqq\{1,\dots,J\}$. Each expert $j\in[J]$ is associated with a measurable cost function $c_{j}:\mathcal{X}\times\mathcal{Y}\to[0,C]$, that quantifies the expense of routing $(x,y)$ to expert $j$. This formulation is agnostic to the nature of the expert: an expert may be a predictive model, a language model system, a human annotator, or any other decision mechanism that induces measurable costs. We model expert prediction as $e_{j}:\mathcal{X}\to\mathcal{Y}$ and decompose the cost as $c_{j}(x,y)\;\coloneqq\;\psi\bigl(e_{j}(x),\,y\bigr)+\beta_{j}$, where $\psi:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a measurable task-specific loss (e.g., the 011 loss for classification) and $\beta_{j}\geq 0$ is an expert consultation cost (Madras et al., 2018; Mozannar and Sontag, 2020; Verma et al., 2022). We denote the tuple of expert predictors by $\mathbf{e}\coloneqq(e_{1},\dots,e_{J})$.
Given a collection of fixed, pre-trained experts, L2D seeks a routing policy that assigns each instance to the most cost-effective expert (Madras et al., 2018; Mozannar and Sontag, 2020; Narasimhan et al., 2022; Mao et al., 2023a, 2024c; Montreuil et al., 2025b, 2026a). Let $\mathcal{H}_{r}\subseteq\{s_{r}:\mathcal{X}\times[J]\to\mathbb{R}\}$ denote a hypothesis class of measurable scoring functions for routing. Each $s_{r}\in\mathcal{H}_{r}$ induces a measurable router $r:\mathcal{X}\to[J]$ via $r(x)=\operatorname*{arg\,max}_{j\in[J]}s_{r}(x,j)$; we write $\mathbf{s}_{r}(x)=(s_{r}(x,j))_{j\in[J]}$ for the full score vector. The deferral loss incurred by routing a realization $(x,y)$ through $r$ is
@@
The indicator functions in $\ell_{\mathrm{def}}$ make it discontinuous and non-differentiable, precluding direct gradient-based optimization (Zhang, 2002; Steinwart, 2007; Mozannar and Sontag, 2020). Standard practice replaces it with a convex surrogate $\Phi_{\mathrm{def}}$ that operates on the scoring function $s_{r}$ and is amenable to gradient-based methods. The surrogate deferral risk is $\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})=\mathbb{E}_{(X,Y)\sim\mathcal{D}}[\Phi_{\mathrm{def}}(s_{r};\,X,Y,\mathbf{e})]$, with best-in-class value $\mathcal{E}_{\Phi_{\mathrm{def}}}^{\ast}(\mathcal{H}_{r})\coloneqq\inf_{s_{r}\in\mathcal{H}_{r}}\mathcal{E}_{\Phi_{\mathrm{def}}}(s_{r})$. To obtain population-level excess-risk control, Awasthi et al. (2022) introduced $\mathcal{H}$-consistency bounds.
@@
When $\mathcal{H}_{r}$ is rich enough for these gaps to vanish, the bound recovers Bayes consistency (Steinwart, 2007; Awasthi et al., 2022). In the standard L2D setting, such bounds have been established for the comp-sum family of cross-entropy surrogates (Mao et al., 2023b, 2024a), guaranteeing that surrogate minimization over a rich enough class recovers the Bayes-optimal router.
@@ 4.2 The Deferral-Advice Loss
A deferral-advice policy is a pair of measurable maps $r:\mathcal{X}\to[J]$ and $q:\mathcal{X}\times[J]\to[K]_{0}$. Let $\mathcal{H}_{q}\subseteq\{s_{q}:\mathcal{X}\times[J]\times[K]_{0}\to\mathbb{R}\}$ be a hypothesis class of measurable query scoring functions; each $s_{q}\in\mathcal{H}_{q}$ induces a query function via $q(x,j)=\operatorname*{arg\,max}_{k\in[K]_{0}}s_{q}(x,j,k)$. We write $\mathbf{S}_{q}(x)\in\mathbb{R}^{J\times(K+1)}$ for the matrix with entries $s_{q}(x,j,k)$, so that each row corresponds to one expert and the row-wise argmax gives the advice index for that expert. Upon observing $x$, the system routes to expert $j=r(x)$ and selects advice index $k=q(x,j)$; the masked advice $\widetilde{A}^{(k)}$ is then revealed to expert $j$. Although $q(x,\cdot)$ is defined for every expert, only the executed value $q(x,r(x))$ enters the realized loss (see Algorithm 2, Figure 1).
@@
The double indicator selects exactly one expert-advice pair per instance. The loss generalizes (1) from $J$ to $J(K+1)$ actions and depends on the policy only through the executed pair (r(x),q(x,r(x)))(r(x),\,q(x,r(x))). Since $c_{j,0}(x,a,y)=c_{j}(x,y)$ by construction, setting $K=0$ recovers the standard deferral loss $\ell_{\mathrm{def}}$ (Mozannar and Sontag, 2020; Mao et al., 2023a; Montreuil et al., 2026a). A concrete numerical example illustrating the loss is given in Example A.3.
@@ 4.3 Deferral with Advice Dominates Standard Deferral
The proof is given in Appendix B.1. The result reveals a fundamental coupling: the router does not compare experts on their unaided costs $\mathbb{E}[c_{j}(X,Y)\mid X=x]$ as in the standard setting, but on their optimally-advised ones. An expert that would never be selected under standard deferral may become optimal once paired with the right advice. The two decisions—who acts and what they see—cannot be separated.
@@
For every expert $j\in[J]$ and $\mathbb{P}_{X}$-a.e. $x$, if $\mathcal{E}^{\star}_{\ell_{\mathrm{def}}}$ denotes the Bayes risk of the standard deferral setting of Section 3, then
@@ 4.4 The Separated Surrogate Fails
Because the protocol is sequential, the natural first attempt is to learn it sequentially as well: one router score, one query head per expert, and smooth logistic replacements for the hard indicators (Bartlett et al., 2006). But does mirroring the control flow suffice for consistency? We show that a broad family of separated surrogates already fails in the smallest non-trivial case: two experts $(J=2)$ and one binary advice $(K=1)$.
@@
has Bayes-optimal expert $j^{\star}=1$, while the unique minimizer of the pointwise surrogate (9) decodes to expert $2$. In particular, the entire family (9) is not Bayes consistent for the true loss.
@@ 4.5 The Augmented Surrogate
A sequential policy $(r,q)$ induces the composite action $\pi(x)=(r(x),\,q(x,r(x)))$, and every composite action $(j,k)$ can in turn be implemented by a sequential policy with the same realized loss. The composite view simply makes explicit what the loss has depended on all along: the executed expert-advice pair. Proposition 16, proved in Appendix B.5, shows that the sequential and composite formulations therefore induce the same population risks and share the same Bayes-optimal decisions as those identified in Lemma 3.
@@
The composite view aligns the learning problem with the Bayes comparison itself. One still faces the non-differentiability of the true loss from Definition 2, so the remaining task is to construct a surrogate on the augmented action space. The key step, proved in Appendix B.6, is a decomposition of the realized loss into an action-independent offset and a weighted sum of composite mismatch indicators $\mathbf{1}\{\pi(x)\neq(j,k)\}$. Since the offset does not affect the pointwise comparison across actions, surrogate design can focus entirely on the mismatch term.
@@
The augmented surrogate succeeds where the separated one fails because it never profiles out a subset of scores. The separated construction optimizes query scores row by row and hands the router a scalar summary per expert; that summary depends on the non-minimal entry and distorts the comparison (Theorem 6). The augmented surrogate, by contrast, assigns one score to each composite action $(j,k)$ and compares them all at once. No row-by-row reduction ever occurs, so the surrogate comparison operates on the same objects as the Bayes comparison.
We now establish this intuition formally: the augmented surrogate admits a quantitative excess-risk transfer bound for the entire comp-sum family and hence for every $\tau\geq 0$.
@@
Corollary 9 (Bayes consistency).
@@
Corollary 9 is the statistical payoff: whenever the hypothesis class is rich enough and the surrogate is well-optimized, the induced policy converges to the Bayes-optimal deferral-advice risk (Steinwart, 2007) defined in Lemma 3. When $\mathcal{H}_{\pi}$ contains all measurable functions, both minimizability gaps vanish by definition (Zhang and Agarwal, 2020; Bartlett et al., 2006; Long and Servedio, 2013), and the bound reduces to classical Bayes-consistency. The separated surrogate offers no such guarantee—its minimizers can converge to a provably suboptimal policy regardless of model capacity.